OpenStudy (anonymous):

Partial Fraction Integral 1/(121-196x^2) dx

OpenStudy (anonymous):

\[\frac{ 1 }{ 121-196x^2 } = \int\limits_{}^{} \frac{ A }{ 11+14x } + \frac{ B }{ 11-14x }\]

OpenStudy (anonymous):

I then multiply everything by (11+14x)(11-14x).

OpenStudy (anonymous):

\[1 = A(11-14x) + B(11+14x)\]

OpenStudy (anonymous):

Then I rewrite it.

OpenStudy (anonymous):

\[1 = 14(-A+B)x + 11(A+B)\] This is where I am stuck.

zepdrix (zepdrix):

You don't really want to group the terms like that. From this step:\[\large 1 = A(11-14x) + B(11+14x)\] You have a couple of options, You can substitute in \(\large x=\dfrac{11}{14}\) which will allow you to solve for B. After which you can find another value to sub in that will cancel out the B term, to solve for A. If that method is too confusing, you can expand everything out and equate like terms. This method will leave you with a system of equations.

OpenStudy (anonymous):

From Mathematica:\[\frac{1}{121-196 x^2}=\frac{1}{22 (14 x+11)}-\frac{1}{22 (14 x-11)} \]

OpenStudy (anonymous):

So the goal is to keep the letters factored out?

zepdrix (zepdrix):

Umm, I guess we can try the second method if that explanation is a bit confusing. Equating like terms from the step you last did,\[\large 1 = 14(-A+B)x + 11(A+B)\] We can equate the constant values on each side,\[\large \color{royalblue}{1=11(A+B)}\]And we can equate the x^1 values on each side,\[\large 0x=14(-A+B)x\]Dividing through by x simplifies this second equation to,\[\large \color{royalblue}{0=14(-A+B)}\] And from here, you have a system of 2 equations that can easily be solved.

OpenStudy (anonymous):

Still struggling, now they we have equated the right side with the left. How do I solve for A and B?

OpenStudy (anonymous):

Does 11 become A? and B become 14 based on position?

zepdrix (zepdrix):

Based on position? I'm not quite sure what that means. We can solve for A and B either through `substitution` or `elimination`. So we have this system,\[1=11A+11B\]\[0=-14A+14B\] Multiply the first equation by 14/11, giving us,\[\frac{14}{11}=14A+14B\]\[0=-14A+14B\] From here, you can add the two equations together, and the A's will cancel out ~ Allowing you to solve for B.

zepdrix (zepdrix):

This is the elimination method ^

OpenStudy (anonymous):

Ahh OK, I get it now! I was trying to each one independently and not as a system. That explains so much! Thank you!

zepdrix (zepdrix):

Ah I see :)